How many gallons of water should be added to 30 gallons of a solution that is 70% antifreeze in order to obtain a mixture that is 60% antifreeze?
Answer by richwmiller(17219)


from the standpoint of antifreeze
.7*30=.60*(30+x)
2.10=1.80+.6x
.30=.6x
5=x
x= 5 gallons of water
from the standpoint of water
(.3*30)+x=(30+x)*.4
9+x=12.+.4x
.6x=3
multiply by 10
6x=30
x=5 gallons of water